OpenStudy (anonymous):

What is the solution to the rational equation (5/(x^2+3x+2))+(1/(x+2))=(1/(3x+3))

OpenStudy (anonymous):

x = -8

hero (hero):

manuel1234, you have given an answer without explanation. That is against CoC. http://openstudy.com/code-of-conduct

hero (hero):

I have to report you

OpenStudy (anonymous):

nooooo dont

hero (hero):

The purpose of this site is to help users. Giving answers does not help anyone learn anything. Giving answers is a form of cheating and cheating is not what this site is about.

OpenStudy (anonymous):

...

hero (hero):

@militarygirl96 , you should want explanations, not answers.

OpenStudy (anonymous):

i want both but this is one that is trial and error so that would be alot of work to show

hero (hero):

It's not trial and error. The steps are simple

hero (hero):

You just have to make sure the denominators are factored first

hero (hero):

\[\huge\color{salmon}{\text{Zepp}}\]

hero (hero):

@zepp

OpenStudy (zepp):

o.o

OpenStudy (zepp):

\[\large \frac{5}{(x^2+3x+2}+\frac{1}{x+2}=\frac{1}{(3x+3)}\]

hero (hero):

\( \humongous \color{orange} {\text{ZEPP}} \) \( \giant \color{orange} {\text{ZEPP}} \) \( \enormous \color{orange} {\text{ZEPP}} \) \( \gigantic \color{orange} {\text{ZEPP}} \) \( \xlarge \color{orange} {\text{ZEPP}} \) \( \enlarged \color{orange} {\text{ZEPP}} \) none of these work

hero (hero):

\[\big\color{red}{\text{ZEPP}}\]

hero (hero):

I don't know why that isn't working

OpenStudy (zepp):

First step would be to get the same denominator\[\large \begin{align} \frac{5}{x^2+3x+2}+\frac{1}{x+2}&=\frac{1}{(3x+3)}\\ \\ \frac{5}{(x+1)(x+2)}+\frac{1}{x+2}&=\frac{1}{(3x+3)}\\ \\ \frac{5}{(x+1)(x+2)}+\frac{1(x+1)}{(x+2)(x+1)}&=\frac{1}{(3x+3)}\\ \\ \\ \frac{5+1(x+1)}{(x+2)(x+1)}&=\frac{1}{3*(x+1)}\\ \\ \\ \frac{5+x+1}{(x+2)(x+1)}&=\frac{1}{3*(x+1)}\\ \\ \frac{x+6}{(x+2)(x+1)}&=\frac{1}{3*(x+1)}\\ \\ \\ (x+2)(x+1)&=(x+6)*3*(x+1)\\ \\ \frac{(x+2)(x+1)}{(x+1)}&=3(x+6)\\ \\ \\ x+2&=3x+18\\ \\ x&=3x+18-2\\ \\ x&=3x+16\\ \\ x-3x&=16\\ \\ -2x&=16\\ \\ x&=\frac{16}{-2}\\ \\ x&=-8 \end{align}\]

hero (hero):

My steps would be better

OpenStudy (zepp):

What's with all these adjective in front of my name, @Hero? o.O

hero (hero):

Fails at getting bigger font size

OpenStudy (zepp):

Post your steps please? :D

OpenStudy (zepp):

I skipped the first factorization thing, otherwise it would be longer :S

hero (hero):

Nah, that would be a complete waste of time

hero (hero):

I can show you in vyew though

OpenStudy (zepp):

Hero you lazy ;C

OpenStudy (zepp):

Nah it's fine, I'm watching a lecture, thanks for the proposition though

hero (hero):

Basically I can do these without using LCD